# A Question-centric Multi-experts Contrastive Learning Framework for Improving the Accuracy and Interpretability of Deep Sequential Knowledge Tracing Models

### v1

@@ Abstract.
Knowledge tracing (KT) plays a crucial role in predicting students’ future performance by analyzing their historical learning processes. Deep neural networks (DNNs) have shown great potential in solving the KT problem. However, there still exist some important challenges when applying deep learning techniques to model the KT process. The first challenge lies in taking the individual information of the question into modeling. This is crucial because, despite questions sharing the same knowledge component (KC) [^1] A KC is a generalization of everyday terms like concept, skill, or principle. In the KT task, questions are often associated with one or more KCs, which represent the underlying knowledge that a student needs to grasp in order to answer the question correctly. In this paper, we directly regard “concept” as KC., students’ knowledge acquisition on homogeneous questions [^2] In this paper, we refer to questions that share the same set of KCs as “homogeneous questions”. can vary significantly. However, due to the large question bank, the average number of interactions per question may not be sufficient [^3] In contrast, the average number of interactions per KC is much larger than per question, because the KC bank is much smaller than the question bank, more details are illustrated in Table 1.. This limitation can potentially result in overfitting of the question embedding and inaccurate question knowledge acquisition state [^4] Question knowledge acquisition state denotes the student’s knowledge acquisition level for each question within the question bank, it is obtained by the corresponding question embedding. More details are illustrated in Section 4.2. that relies on its corresponding question representation. Furthermore, there is a considerable portion of questions [^5] In this paper, we regard these questions as “less interactive questions”. receiving relatively less interaction from students in comparison to the majority of questions. This can further increase the risk of overfitting and lower the accuracy of the obtained question knowledge acquisition state. The second challenge lies in interpreting the prediction results from existing deep learning-based KT models. In real-world applications, while it may not be necessary to have complete transparency and interpretability of the model parameters, it is crucial to present the model’s prediction results in a manner that teachers find interpretable. This makes teachers accept the rationale behind the prediction results and utilize them to design teaching activities and tailored learning strategies for students. However, the inherent black-box nature of deep learning techniques often poses a hurdle for teachers to fully embrace the model’s prediction results. To address these challenges, we propose a Question-centric Multi-experts Contrastive Learning framework for KT called Q-MCKT. This framework explicitly models students’ knowledge acquisition state at both the question and concept levels. It leverages the mixture of experts technique to capture a more robust and accurate knowledge acquisition state in both question and concept levels for prediction. Additionally, a fine-grained question-centric contrastive task is introduced to enhance the representations of less interactive questions and improve the accuracy of their corresponding question knowledge acquisition states. Moreover, Q-MCKT utilizes an item response theory-based prediction layer to generate interpretable prediction results based on the knowledge acquisition states obtained from the question and concept knowledge acquisition modules. We evaluate the proposed Q-MCKT framework on four public real-world educational datasets. The experimental results demonstrate that our approach outperforms a wide range of deep learning-based KT models in terms of prediction accuracy while maintaining better model interpretability. To ensure reproducibility, we have provided all the datasets and code on our website at https://pykt.org/.
@@
copyright: noneccs: Social and professional topics Student assessmentccs: Applied computing Learning management systems
@@ 1. Introduction
In recent years, the utilization of online educational platforms, such as massive open online courses (MOOCs) [^6] https://www.mooc.org and Coursera [^7] https://www.coursera.org, has significantly increased, particularly due to the impact of the COVID-19 pandemic and subsequent school closures. The surge in online learning has resulted in a wealth of student learning data. Furthermore, recent advancements in data analytics have greatly contributed to significant progress in the collection and analysis of extensive learner data. The availability of vast amounts of learner data, coupled with sophisticated data analytics, has opened up new possibilities for leveraging artificial intelligence (AI) in education settings. Since every student is unique and has distinct learning needs, personalized learning has emerged as a crucial aspect of education. As a result, the development of personalized online education platforms, such as intelligent tutoring systems (ITS) (Woolf, 2010), has become a prominent focus within the field of AI in education. The ability to assess each student’s current knowledge level and automatically provide personalized feedback (Piech et al., 2015b) and recommendations (Lan and Baraniuk, 2016) to the students through the large-scale learning data obtained from the online learning environment is the key to the success of ITS.
@@
Knowledge tracing is an important task in the field of Intelligent Tutoring Systems (ITS) (Lee et al., 2022; Sonkar et al., 2020; Chen et al., 2023). It aims to predict the outcomes of students over questions by modeling their mastery of knowledge, i.e., knowledge states, as they interact with learning platforms. Take Figure 1 as an example, student $s_{1}$ has sequentially answered five questions ($q_{1}$~$q_{5}$) where $q_{1}$ and $q_{5}$ are answered correctly while $q_{2}$ and $q_{4}$ are answered incorrectly. This indicates that $s_{1}$ may have a good mastery of the “Integer Distributive Law” and “Multi-digit Calculation” KCs, but is not familiar with the “Mixed Basic Operations” and “Factor in Integers” KCs1 (the KCs associated with each question can be found in the Question-KC Relation Matrix in Figure 1). Based on the current mastery of knowledge, the KT task can predict the performance of student in the following sixth question $q_{6}$. Once the knowledge mastery is acquired through KT, students can find out their weaknesses in time and then carry out targeted exercises. It can also aid online learning platforms in recommending personalized learning materials to cater to the unique needs of individual students (Wu et al., 2020; Liu et al., 2019).
@@
The KT related research has been studied since the 1990s when Corbett and Anderson (1994a), to the best of our knowledge, were the first to estimate students’ current knowledge with regard to each individual KC. Since then, many attempts have been made to solve the KT problem, such as probabilistic graphical models (Käser et al., 2017; Hawkins et al., 2014) and factor analysis-based models (Cen et al., 2006; Lavoué et al., 2018; Thai-Nghe et al., 2012). While these approaches achieve encouraging progress, they always rely on simplified assumptions such as no forgetting during the learning process (Corbett and Anderson, 1994a), limiting their applications in real scenarios. Moreover, these methods are always hand-crafted engineerings. However, incorporating all potential factors into the KT model is essentially a non-trival task.
Figure 2. The distribution of question interaction frequency in two real-world datasets is depicted in the figure. It is evident that in the NIPS34 dataset, over 25% of questions have an interaction frequency of less than 300, which is comparatively lower than other questions. Similarly, in the EdNet dataset, more than 31% of questions are associated with an interaction frequency of less than 20.
@@
Recently, with the rapid development of deep neural networks, many deep learning based knowledge tracing (DLKT) models are developed (Chen et al., 2023, 2018; Guo et al., 2021; Long et al., 2021a; Piech et al., 2015a; Yeung and Yeung, 2018a; Abdelrahman and Wang, 2019; Yeung, 2019a; Zhang et al., 2017; Choi et al., 2020a; Ghosh et al., 2020; Pandey and Karypis, 2019; Pandey and Srivastava, 2020a). One of the representative approaches among them is the deep sequential KT modeling, which utilizes auto-regressive architectures, such as LSTM and GRU (Hochreiter and Schmidhuber, 1997; Chung et al., 2014), to represent student’s knowledge states (e.g., the mastery level of KCs) as the hidden states of recurrent units (Chen et al., 2023, 2018; Guo et al., 2021; Long et al., 2021a; Piech et al., 2015a; Yeung and Yeung, 2018a). Due to the ability to learn sequential dependencies from student interaction data, deep sequential KT models draw attention from researchers from different communities and achieve great success in improving the KT prediction accuracy (Minn et al., 2018; Nagatani et al., 2019a; Su et al., 2018). In spite of the promising results demonstrated by previous methods, some important challenges still exist when applying deep sequential KT models on real-world educational data. Firstly, taking the individual information of the question into modeling is challenging for the KT model. This challenge is caused by the large question bank, which leads to an average number of interactions per question may not be sufficient. Thus, it can potentially result in overfitting of the question embedding and inaccurate question knowledge acquisition state4 that relies on their corresponding question representation. Additionally, there exists a considerable portion of less interactive questions5 in the dataset, which have a higher risk of getting overfitted. For example, as shown in Figure 2, more than 31% of questions are associated with an interaction frequency of less than 20 in the EdNet dataset, and 25% of questions have an interaction frequency of less than 300 in the NIPS34 dataset. Consequently, it is difficult for KT model to accurately model the question knowledge acquisition state, especially when encountering the less interactive question. Secondly, despite the significant improvements in prediction accuracy achieved by deep learning-based knowledge tracing (DLKT) models compared to traditional cognitive models, there is still a challenge in extracting psychologically interpretable prediction results from them. This lack of interpretability often leaves teachers dissatisfied, as they require interpretable prediction results before assigning appropriate learning materials for students (although the model parameters are not explainable, the rationale behind the prediction results should be interpretable for teachers to accept).
@@
In this paper, we address aforementioned challenges by proposing a novel framework called Question-centric Multi-experts Contrastive Learning framework for KT, i.e., Q-MCKT. Specifically, the question knowledge acquisition module of Q-MCKT estimates the question knowledge acquisition state4 of students by modeling the variations in students’ question knowledge states after receiving responses to specific questions. To mitigate the potential inaccuracy of the question knowledge acquisition state caused by the overfitted question embedding, we draw inspiration from the realistic educational scenario where multiple teachers evaluate a student’s performance to obtain an overall score (as illustrated in Figure 3) and incorporate the mixture of experts technique (Jacobs et al., 1991) into Q-MCKT to obtain a more robust and accurate question knowledge acquisition state. For less interactive questions, Q-MCKT further employs an auxiliary fine-grained question-centric contrastive learning procedure to enhance its question representations and thereby improves the accuracy of its corresponding question knowledge acquisition state. Furthermore, in real educational settings, students often make guesses or slip-ups when interacting with certain questions, which can lead the question knowledge acquisition state overly sensitive to individual interactions. Therefore, Q-MCKT also incorporates a concept knowledge acquisition module in a similar manner to the question knowledge acquisition module. This module aims to obtain the concept knowledge acquisition state [^8] Concept knowledge acquisition state denotes the student’s acquisition level for each concept. It is obtained by the concept embedding, more details are illustrated in Section 4.3, which serves as the supplemental information for the question knowledge acquisition state. To enhance the interpretability of prediction results, Q-MCKT directly feeds the knowledge acquisition states from question and concept knowledge acquisition modules into an interpretable prediction layer. This interpretable prediction layer utilizes the Item Response Theory (IRT) (Rasch, 1993) from psychometrics, which is widely used in educational scenarios. We evaluate Q-MCKT on four benchmark datasets by comparing it with 15 previous approaches under a rigorous KT evaluation protocol (Liu et al., 2022). The experimental results demonstrate that Q-MCKT achieves superior prediction performance and psychologically meaningful interpretability simultaneously.
@@ 2.1. Knowledge Tracing
There exist two paradigms of knowledge tracing techniques: conventional methods and deep learning-based methods. When it comes to conventional methods, one of the most renowned methods is Bayesian knowledge tracing (BKT) (Corbett and Anderson, 1994b; Yudelson et al., 2013). BKT utilizes binary variables to depict the knowledge state of students and leverages the Hidden Markov Model to capture the sequential progression of knowledge states. Subsequent studies have delved into integrating additional factors (e.g. question difficulty (Pardos and Heffernan, 2011), guess estimation (Baker et al., 2008) and student ID (Vie and Kashima, 2019)) related to learning scenarios into the modeling process. These investigations have revealed that as the number of incorporated modeling factors grows, the performance of the models improves correspondingly. As deep learning techniques have achieved great success in various applications, it has also been applied in KT problem to model student’s historical learning interactions and enable a comprehensive understanding of the dependencies within the learning interactions, leading to enhanced predictive capabilities. The recent deep learning-based methods can be categorized as follows:
@@
  • Deep sequential models: the chronologically ordered interaction sequence is captured by deep sequential models such as LSTM and GRU (Chen et al., 2018; Guo et al., 2021; Lee and Yeung, 2019; Liu et al., 2019; Minn et al., 2018; Nagatani et al., 2019a; Piech et al., 2015a; Su et al., 2018; Yeung and Yeung, 2018b). For example, Piech et al. (2015a) propose deep knowledge tracing (DKT) model, which employs the hidden state of LSTM to represent student’s knowledge state and then feeds the hidden state into a binary classifier to predict the students’ response performances. Yeung and Yeung (2018a) utilize prediction-consistent regularization technique to address the problems of input reconstruction failure and prediction inconsistency in DKT (Piech et al., 2015a). Motivated by the learning curve theory (Murray et al., 2013)Nagatani et al. (2019b) take student’s forgetting behavior into consideration to enhance DKT (Piech et al., 2015a). Lee and Yeung (2019) use student knowledge state encoder and skill encoder to predict the student response performance via the dot product.
@@
  • Memory augmented models: the latent relationships among KCs are effectively captured through the utilization of an external memory structure, especially key-value memory, which is updated iteratively (Abdelrahman and Wang, 2019; Shen et al., 2021; Zhang et al., 2017). For example, Zhang et al. (2017) incorporate a well-designed static key matrix to store the interconnections between various KCs, while simultaneously employing a dynamic value matrix to iteratively update the students’ knowledge state.
@@
  • Adversarial based models: the adversarial training techniques such as adversarial perturbations are applied into the original student interaction sequence to reduce the risk of DLKT overfitting and limited generalization problem. For example, Guo et al. (2021) enhance the generalization ability of the DLKT model by introducing adversarial perturbations into the student interaction sequence at embedding level. Due to its carefully crafted perturbations into the sequence, the augmenting model can generalize well to diverse student interactions.
@@
  • Graph based models: the response interactions between students and questions and the knowledge associations between questions and KCs form a tripartite graph and graph based techniques are applied to aggregate such relations (Nakagawa et al., 2019; Tong et al., 2020; Yang et al., 2020). For example, Nakagawa et al. (2019) utilize graph neural network to aggregate the node features related to the answered concept and subsequently updates the student’s knowledge states.
@@
  • Attention based models: dependence between interactions is captured by the attention mechanism and its variants (Ghosh et al., 2020; Pandey and Srivastava, 2020b; Pu et al., 2020; Zhang et al., 2021). For example, Ghosh et al. (2020) leverage an attention mechanism to characterize the time distance between questions and the past interaction of students. Pandey and Karypis (2019) utilize a self-attention mechanism to capture relations between exercises and the student responses. Pandey and Srivastava (2020b) introduce a relation-aware self-attention layer that takes the individual question textual content and student past performance to rich the contextual information of student learning process. Choi et al. (2020a) use the Transformer-based encoder-decoder architecture to capture students’ exercise and response sequences.

### v2

@@ Abstract.
Knowledge tracing (KT) plays a crucial role in predicting students’ future performance by analyzing their historical learning processes. Deep neural networks (DNNs) have shown great potential in solving the KT problem. However, there still exist some important challenges when applying deep learning techniques to model the KT process. The first challenge lies in modeling the individual question information. This is crucial because students’ knowledge acquisition on questions that share the same set of knowledge components (KCs) may vary significantly. However, due to the large question bank, the average number of interactions per question may not be sufficient. This limitation can potentially result in overfitting of the question embedding and inaccurate question knowledge acquisition state that relies on its corresponding question representation. Furthermore, there is a considerable portion of questions receiving relatively less interaction from students in comparison to the majority of questions. This can further increase the risk of overfitting and lower the accuracy of the obtained question knowledge acquisition state. The second challenge lies in interpreting the prediction results from existing deep learning-based KT models. In real-world applications, while it may not be necessary to have complete transparency and interpretability of the model parameters, it is crucial to present the model’s prediction results in a manner that teachers find interpretable. This makes teachers accept the rationale behind the prediction results and utilize them to design teaching activities and tailored learning strategies for students. However, the inherent black-box nature of deep learning techniques often poses a hurdle for teachers to fully embrace the model’s prediction results. To address these challenges, we propose a Question-centric Multi-experts Contrastive Learning framework for KT called Q-MCKT. This framework explicitly models students’ knowledge acquisition state at both the question and concept levels. It leverages the mixture of experts technique to capture a more robust and accurate knowledge acquisition state in both question and concept levels for prediction. Additionally, a fine-grained question-centric contrastive learning task is introduced to enhance the representations of less interactive questions and improve the accuracy of their corresponding question knowledge acquisition states. Moreover, Q-MCKT utilizes an item response theory-based prediction layer to generate interpretable prediction results based on the knowledge acquisition states obtained from the question and concept knowledge acquisition modules. We evaluate the proposed Q-MCKT framework on four public real-world educational datasets. The experimental results demonstrate that our approach outperforms a wide range of deep learning-based KT models in terms of prediction accuracy while maintaining better model interpretability. To ensure reproducibility, we have provided all the datasets and code on our website at https://github.com/rattlesnakey/Q-MCKT.
@@
copyright: nonedoi: ccs: Applied computing Educationccs: Computing methodologies Machine learning
@@ 1. Introduction
In recent years, the utilization of online educational platforms, such as massive open online courses (MOOCs) and Coursera, has significantly increased, particularly due to the impact of the COVID-19 pandemic and subsequent school closures. The surge in online learning has resulted in a wealth of student learning data. Furthermore, recent advancements in data analytics have greatly contributed to significant progress in the collection and analysis of extensive learner data. The availability of vast amounts of learner data, coupled with sophisticated data analytics, has opened up new possibilities for leveraging AI in education settings. Since every student is unique and has distinct learning needs, personalized learning has emerged as a crucial aspect of education. As a result, the development of personalized online education platforms, such as intelligent tutoring systems (ITS) (Woolf, 2010), has become a prominent focus within the field of AI in education. The ability to assess each student’s current knowledge level and automatically provide personalized feedback (Piech et al., 2015b) and recommendations (Lan and Baraniuk, 2016) to the students through the large-scale learning data obtained from the online learning environment is the key to the success of ITS.
@@
Knowledge tracing is an important task in the field of ITS (Lee et al., 2022; Sonkar et al., 2020; Chen et al., 2023). It aims to predict the outcomes of students over questions by modeling their mastery of knowledge, i.e., knowledge states, as they interact with learning platforms. Take Figure 1 as an example, student $s_{1}$ has sequentially answered five questions ($q_{1}$~$q_{5}$) where $q_{1}$ and $q_{5}$ are answered correctly while $q_{2}$ and $q_{4}$ are answered incorrectly. This indicates that $s_{1}$ may have a good mastery of the “Integer Distributive Law” and “Multi-digit Calculation” KCs [^1] A KC is a generalization of everyday terms like concept, skill, or principle. In the KT task, questions are often associated with one or more KCs, which represent the underlying knowledge that a student needs to grasp in order to answer the question correctly. In this paper, we directly regard “concept” as KC., but is not familiar with the “Mixed Basic Operations” and “Factor in Integers”. The KCs associated with each question can be found in the Question-KC Relation Matrix in Figure 1. Based on the current mastery of knowledge, the KT task can predict the performance of student in the following sixth question $q_{6}$. Once the knowledge mastery is acquired through KT, students can find out their weaknesses in time and then carry out targeted exercises. It can also aid online learning platforms in recommending personalized learning materials to cater to the unique needs of individual students (Wu et al., 2020; Liu et al., 2019).
@@
The KT related research has been studied since the 1990s when Corbett and Anderson (1994), to the best of our knowledge, were the first to estimate students’ current knowledge with regard to each individual KC. Since then, many attempts have been made to solve the KT problem, such as probabilistic graphical models (Käser et al., 2017; Hawkins et al., 2014) and factor analysis-based models (Cen et al., 2006; Lavoué et al., 2018; Thai-Nghe et al., 2012). While these approaches achieve encouraging progress, they always rely on simplified assumptions such as no forgetting during the learning process (Corbett and Anderson, 1994), limiting their applications in real scenarios. Moreover, these methods are always hand-crafted engineerings. However, incorporating all potential factors into the KT model is essentially a non-trivial task.
Figure 2. The distribution of question interaction frequency in two real-world datasets.
@@
Recently, with the rapid development of deep neural networks, many DLKT models are developed (Chen et al., 2023, 2018; Guo et al., 2021; Long et al., 2021; Piech et al., 2015a; Yeung and Yeung, 2018; Abdelrahman and Wang, 2019; Yeung, 2019a; Zhang et al., 2017; Choi et al., 2020a; Ghosh et al., 2020; Pandey and Karypis, 2019; Pandey and Srivastava, 2020). One of the representative approaches among them is the deep sequential KT modeling, which utilizes auto-regressive architectures, such as LSTM and GRU (Hochreiter and Schmidhuber, 1997; Chung et al., 2014), to represent student’s knowledge states (e.g., the mastery level of KCs) as the hidden states of recurrent units (Chen et al., 2023, 2018; Guo et al., 2021; Long et al., 2021; Piech et al., 2015a; Yeung and Yeung, 2018). Due to the ability to learn sequential dependencies from student interaction data, deep sequential KT models draw attention from researchers from different communities and achieve great success in improving the KT prediction accuracy (Minn et al., 2018; Nagatani et al., 2019; Su et al., 2018). In spite of the promising results demonstrated by previous methods, some important challenges still exist when applying deep sequential KT models on real-world educational data. Firstly, taking the individual information of the question into modeling is challenging for the KT model. This challenge is caused by the large question bank, which leads to an average number of interactions per question may not be sufficient. Thus, it can potentially result in overfitting of the question embedding and inaccurate question knowledge acquisition state [^2] Question knowledge acquisition state denotes the student’s knowledge acquisition level for each question within the question bank, it is obtained by the corresponding question embedding. More details are illustrated in Section 4.2. that relies on their corresponding question representation. Additionally, there exists a considerable portion of less interactive questions in the dataset, which have a higher risk of getting overfitted. For example, as shown in Figure 2, more than 31% of questions are associated with an interaction frequency of less than 20 in the EdNet dataset, and 25% of questions have an interaction frequency of less than 300 in the NIPS34 dataset. Consequently, it is difficult for KT model to accurately model the question knowledge acquisition state, especially when encountering the less interactive question. Secondly, despite the significant improvements in prediction accuracy achieved by DLKT models compared to traditional cognitive models, there is still a challenge in extracting psychologically interpretable prediction results from them. This lack of interpretability often leaves teachers dissatisfied, as they require interpretable prediction results before assigning appropriate learning materials for students. It should be noted that although the model parameters may not be explainable, the reasoning behind the prediction results must be interpretable for teachers to accept.
@@
In this paper, we address aforementioned challenges by proposing a novel framework called Question-centric Multi-experts Contrastive Learning framework for KT, i.e., Q-MCKT. Specifically, the question knowledge acquisition module of Q-MCKT estimates the question knowledge acquisition state2 of students by modeling the variations in students’ question knowledge states after receiving responses to specific questions. To mitigate the potential inaccuracy of the question knowledge acquisition state caused by the overfitted question embedding, we draw inspiration from the realistic educational scenario where multiple teachers evaluate a student’s performance to obtain an overall score (as illustrated in Figure 3) and incorporate the mixture of experts technique (Jacobs et al., 1991) into Q-MCKT to obtain a more robust and accurate question knowledge acquisition state. For less interactive questions, Q-MCKT further employs an auxiliary fine-grained question-centric contrastive learning procedure to enhance its question representations and thereby improves the accuracy of its corresponding question knowledge acquisition state. Furthermore, in real educational settings, students often make guesses or slip-ups when interacting with certain questions, which can lead the question knowledge acquisition state overly sensitive to individual interactions. Therefore, Q-MCKT also incorporates a concept knowledge acquisition module in a similar manner to the question knowledge acquisition module. This module aims to obtain the concept knowledge acquisition state [^3] Concept knowledge acquisition state denotes the student’s acquisition level for each concept. It is obtained by the concept embedding, more details are illustrated in Section 4.3., which serves as the supplemental information for the question knowledge acquisition state. To enhance the interpretability of prediction results, Q-MCKT directly feeds the knowledge acquisition states from question and concept knowledge acquisition modules into an interpretable prediction layer. This interpretable prediction layer utilizes the Item Response Theory (IRT) (Rasch, 1993) from psychometrics, which is widely used in educational scenarios. We evaluate Q-MCKT on four benchmark datasets by comparing it with 15 previous approaches under a rigorous KT evaluation protocol (Liu et al., 2022). The experimental results demonstrate that Q-MCKT achieves superior prediction performance and psychologically meaningful interpretability simultaneously.
@@ 2.1. Knowledge Tracing
There exist two paradigms of knowledge tracing techniques: conventional methods and deep learning-based methods. When it comes to conventional methods, one of the most renowned methods is Bayesian knowledge tracing (BKT) (Yudelson et al., 2013). BKT utilized binary variables to depict the knowledge state of students and leveraged the Hidden Markov Model to capture the sequential progression of knowledge states. Subsequent studies have delved into integrating additional factors (e.g. question difficulty (Pardos and Heffernan, 2011), guess estimation (Baker et al., 2008) and student ID (Vie and Kashima, 2019)) related to learning scenarios into the modeling process. These investigations have revealed that as the number of incorporated modeling factors grows, the performance of the models improves correspondingly. As deep learning techniques have achieved great success in various applications, it has also been applied in KT problem to model student’s historical learning interactions and enable a comprehensive understanding of the dependencies within the learning interactions, leading to enhanced predictive capabilities. The recent deep learning-based methods can be categorized as follows:
@@
  • Deep sequential models: the chronologically ordered interaction sequence is captured by deep sequential models such as LSTM and GRU (Chen et al., 2018; Guo et al., 2021; Lee and Yeung, 2019; Liu et al., 2019; Minn et al., 2018; Nagatani et al., 2019; Piech et al., 2015a; Su et al., 2018; Yeung and Yeung, 2018). For example, Piech et al. (2015a) proposed deep knowledge tracing (DKT) model, which employed the hidden state of LSTM to represent student’s knowledge state and then fed the hidden state into a binary classifier to predict the students’ response performances. Yeung and Yeung (2018) utilized prediction-consistent regularization technique to address the problems of input reconstruction failure and prediction inconsistency in DKT (Piech et al., 2015a). Motivated by the learning curve theory (Murray et al., 2013)Nagatani et al. (2019) taked student’s forgetting behavior into consideration to enhance DKT (Piech et al., 2015a). Lee and Yeung (2019) used student knowledge state encoder and skill encoder to predict the student response performance via the dot product.
@@
  • Memory augmented models: the latent relationships among KCs are effectively captured through the utilization of an external memory structure, especially key-value memory, which is updated iteratively (Abdelrahman and Wang, 2019; Shen et al., 2021; Zhang et al., 2017). For example, Zhang et al. (2017) incorporated a well-designed static key matrix to store the interconnections between various KCs, while simultaneously employing a dynamic value matrix to iteratively update the students’ knowledge state.
@@
  • Adversarial based models: the adversarial training techniques such as adversarial perturbations are applied into the original student interaction sequence to reduce the risk of DLKT overfitting and limited generalization problem. For example, Guo et al. (2021) enhanced the generalization ability of the DLKT model by introducing adversarial perturbations into the student interaction sequence at embedding level. Due to its carefully crafted perturbations into the sequence, the augmenting model can generalize well to diverse student interactions.
@@
  • Graph based models: the response interactions between students and questions and the knowledge associations between questions and KCs form a tripartite graph and graph based techniques are applied to aggregate such relations (Nakagawa et al., 2019; Tong et al., 2020; Yang et al., 2020). For example, Nakagawa et al. (2019) utilized graph neural network to aggregate the node features related to the answered concept and subsequently updated the student’s knowledge states.
@@
  • Attention based models: dependence between interactions is captured by the attention mechanism and its variants (Ghosh et al., 2020; Pandey and Srivastava, 2020; Pu et al., 2020; Zhang et al., 2021). For example, Ghosh et al. (2020) leveraged an attention mechanism to characterize the time distance between questions and the past interaction of students. Pandey and Karypis (2019) utilized a self-attention mechanism to capture relations between exercises and the student responses. Pandey and Srivastava (2020) introduced a relation-aware self-attention layer that taked the individual question textual content and student past performance to rich the contextual information of student learning process. Choi et al. (2020a) used the Transformer-based encoder-decoder architecture to capture students’ exercise and response sequences.